OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition fifty point five divided by zero point five A) one hundred one twenty−six point nine minus eleven point six B) fifteen point three one hundred seventy times zero point nine C) ten point one three point nine plus six point two D) one hundred fifty−three

OpenStudy (anonymous):

help!!!!!!

OpenStudy (anonymous):

1. 50.5/0.5 = 101 (A) 2. 26.9-11.6 = 15.3 (B) 3. 170 x 0.9 = 153 (D) 4. 3.9 + 6.2 =10.1 (C)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome